OBSERVATION AND AB INITIO CALCULATION OF THE TWO-PHOTON VIBRONIC SPECTROSCOPY OF JET-COOLED ALLENE AND PROPYNE

JAU-CHIN SHIEH, JIA-LIN CHANG, YI-JI LIN, A. M. MEBEL AND \underline{YIT-TSONG CHEN}, Department of Chemistry, National Taiwan University, Taipei 106, Taiwan, and Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 106, Taiwan.

The two-photon allowed vibronic spectra of jet-cooled allene (H_{2}CCCH_{2}) and propyne (H_{3}CCCH) at 7.5-9 eV have been observed by 2 + 1 REMPI spectroscopy. Excited electronic states of 2^{1}A_{1} (π-3p/π-π^{∗}), 2^{1}A_{2} (π-3p), 2^{1}B_{1} (π-3p) and 3^{1}A_{1} (π-3d/π-π^{∗}) in allene and ^{1}A_{2} (π-π^{∗}) in propyne are responsible for the observed spectra according to recent ab initio calculations. Geometries of the singlet excited electronic states of allene and propyne have been optimized at the CASSCF level of theory with the 6-311(2+)G^{∗∗} basis set. Vertical and adiabatic excitation energies, calculated by the MRCI and EOM-CCSD methods are compared with experiment. Vibrational frequencies and normal coordinates for the ground and excited states are used for the calculations of vibrational overlap integrals and Franck-Condon factors, taking into account distortion, displacement and normal mode mixing (Duschinsky matrix), as we demonstrated previously in the ethylene case.^{a}^{,}^{b} Major features of the observed vibronic spectra of allene and propyne will be interpreted on the basis of computed Franck-Condon factors. REMPI spectra to cover the energy range of 6 eV to ionization potential are undertaken.

aA. M. Mebel, Y.-T. Chen and S. H. Lin, Chem. Phys. Lett., 258, 53 (1996).

bA. M. Mebel, Y.-T. Chen and S. H. Lin, J. Chem. Phys., 105, 9007 (1996).